# Gravitational Wave Memory from Accelerating Relativistic Jets in Multiple Thick Shell Scenarios

### v1

@@ IV.2 Nearby GRB 221009A-like event
An energetic event GRB 221009A occurred nearby (redshift $z=0.151$, corresponding to the luminosity distance of 724 Mpc) [37, 38]. Its bright observed flux in the prompt GRB emission phase reaches Earth every $\sim 10^{4}$ yr. The isotropic equivalent gamma-ray energy of this event was about $E_{{\rm iso,}\gamma}\simeq 1.2\times 10^{55}$ erg. Well-sampled afterglows, spanning from radio to very-high-energy gamma rays, suggest that the event was associated with a narrowly collimated jet characterized by an opening half-angle of $\Delta\theta\sim 0.01$ rad or smaller [39, 40, 41, 42, 43, 44]. Then, the collimation-corrected energy of the gamma-ray emission is estimated as $(\Delta\theta)^{2}E_{{\rm iso,}\gamma}/4\simeq 3\times 10^{50}$ erg. If we assume the efficiency of the prompt GRB emission of 10%, the jet kinetic energy may be $\mathcal{E}\sim 3\times 10^{51}~{}\mathrm{erg}$.
@@
If similar events to GRB 221009A occurred very nearby, then the GW memory is detectable. Because the light curve of GRB 221009A consists of roughly five pulses in $\sim 510$ seconds [38], we adopt $N=5$ and total time $T_{\mathrm{CE}}\sim 510$ s. The distance to the source is assumed to be approximately one-twentieth that of GRB 221009A, i.e., $R=30$ Mpc.
@@
The waveform is plotted in Fig. 14. It shows the detectability to observe the GW memory around 10 Hz and 100 Hz by ET-D. The attenuation at low-frequency begins from $\sim 10^{-3}$ Hz because $fT_{\mathrm{CE}}\geq 1$ with $T_{\mathrm{CE}}^{-1}\sim 10^{-3}$ Hz, which are the exponents of exponential function in Eq. (13). The attenuation due to the shell thickness should appear at the frequency $\delta t_{j}^{-1}\sim(W_{j}/c)^{-1}\sim 10^{4}$ Hz, however, those amplitudes are already below the sensitivity curve of ET-D. This is because the attenuation associated with $T_{\mathrm{obs}}^{-1}\sim 10^{2}~{}$z begins earlier than that of $\delta t_{j}$ under the conditions; $\gamma\in[300,600]$, $r_{0}=10^{7}~{}\mathrm{cm}$, and $\theta\in[0,0.4]$. The characteristic strain determined by Eq. (16) at low-frequency is $\sim 10^{-23}$, which is observed around $10^{-4}$ Hz in Fig. 14. Furthermore, because the magnitude of characteristic strain is almost the same as that observed in the ET-D around $10^{2}$ Hz and is determined by $R$ and $\mathcal{E}$, it may contribute to the estimation of $\mathcal{E}$ of the GRBs with only ET-D observations.
@@
The maximum detectable distance to the source is found to be approximately $\sim 20$ Mpc for ET-D and $\sim 60$ Mpc for DECIGO. If the distance is about 20 Mpc, our hypothetical event is simultaneously detectable by both detectors. Assuming that GW memory from GRB 221009A-like jets is detectable up to a distance of 60 Mpc, the likelihood of observing an on-axis jet is extremely low: about one event every $\sim 2\times 10^{7}$ yrs. However, due to the anti-beaming effect of GW memory emission, the jet’s GW signal is detectable even for off-axis viewing angles. With a beaming factor of $f_{b}=2\pi/\Delta\Omega\sim 2\times 10^{4}$, the detection rate is boosted to roughly one event every $\sim 10^{3}$ yrs. Although the detection rate of such events may remain low, nearby events with GW memory could provide valuable insights into the jet launching mechanism.
@@ IV.3 Possible GW emission from a black hole formation event in M31
Recently, the disappearance of a massive star in M31 was observed, marking the formation of a black hole [45]. The progenitor star, with an initial mass of approximately $20M_{\odot}$, evolved into a red supergiant before collapsing. At the terminal nuclear burning phase, its mass was estimated at $6.7M_{\odot}$ , and the star’s core collapsed and turned into a black hole with a mass $6.5M_{\odot}$. During its core collapse, a part of the stellar material was ejected. One can expect that some of them will likely generate a relativistic jet. However, the lack of bright optical emission suggested that, even if such a jet was launched, it likely did not escape from the star. If the GW memory had been detected, we would have obtained evidence of the jet in the star.
We simulated GW memory from M31 and discussed its detectability. The distance to the source M31 is relatively close, $R=0.77$ Mpc, and we typically set the parameters such as $T_{\mathrm{CE}}\sim 100$ sec and $\gamma_{j}\in[100,300]$. The hypothetical GW memory from M31 based on this event is shown in Fig. 15. Regarding the solid cyan line, we see that the initial attenuation at $10^{-2}$ Hz, induced by $T_{\mathrm{CE}}^{-1}\sim 10^{-2}$ Hz, and further attenuation around $T_{\mathrm{obs}}^{-1}\sim 10^{3}$ Hz. The thickness impact appears around $\delta t_{j}^{-1}\sim 10^{3}$ Hz, although these amplitudes are under the sensitivity curve of ET-D. The maximum amplitude given by Eq. (16) is $\sim 10^{-23}$, which is almost the same as the signal amplitude in the frequency range observable by ET-D. This indicates that a GW memory observed by ET-D is helpful to understand the total energy $\mathcal{E}$ of GRBs, which is the same situation in Fig. 14. Furthermore, the detection by DECIGO and ET-D depends on the $\mathcal{E}$. The dashed purple line in Fig. 15 shows the waveform under the same conditions as the solid cyan line, except for a smaller $\mathcal{E}$. This indicates that even for nearby GW events, detection remains challenging with next-generation detectors when the energy of the central engine is relatively small, around $10^{50}~{}\mathrm{erg}$.
@@ V Summary and Discussion
The time interval of each shell $\delta\tau_{j}$ with multiple shells affects the waveform shape. When the $\delta\tau_{j}={\rm const.}$, which means $W_{j}={\rm const.}$, the waveform has multiple peaks at specific frequencies due to the vanishing attenuation. If such GW memories are observed, the shells must be emitted periodically. In contrast, the waveform loses such specific peaks when the $\delta\tau_{j}$ exhibits disorder, such as each shell emitted on different interval time $\delta\tau_{j}$.
@@
Furthermore, we have shown that the GW memories from relatively energetic phenomena with typical GRBs parameters would be detectable by multiple GW detectors, ET-D and DECIGO.
After the initial acceleration, the GW amplitude on the low-frequency side is saturated, which is determined by the total energy in the jet. In this sense, the GW memory can serve as a form of calorimetry. However, after the initial acceleration, the jet evolves as it propagates as introduced in Sec. I. The amplitude of the GW memory varies with time, and its detailed analysis may tell us the total jet energy. Once the jet energy is known, we can estimate the efficiency of the prompt gamma-ray emission, whose constraints on the emission mechanism are currently unknown. Note that the order of magnitude of the jet kinetic energy of low-luminosity GRBs has not yet been rigorously determined. A popular model is in the low energy jet, while the off-axis jet model gives larger energy [46, 47, 48, 49].
This paper shows that the GW memory from accelerating jets is only detectable for nearby sources. However, multi-messenger signals other than GW are likely to be expected, if the GW memory signatures are detected. In most cases, the jet may be viewed off-axis. Then, electromagnetic emission is dim in gamma rays, and typical internal shock emission may be seen in the X-ray or ultra-violet/optical band due to the relativistic beaming effect [19, 50, 51]. The prompt GRB emission may not be detected and only off-axis afterglow emission may be observable [52, 53, 54]. If the GRB jet contains baryons to some extent, one can also expect high-energy (TeV–PeV) neutrinos [55, 35, 56, 57, 58]. When the jet is viewed off-axis, the neutrino emission is again dim but may still be observable [59].
@@ Acknowledgements.
We would like to thank Drs. K. Ioka, S. J. Tanaka, S. Kimura and K. Komori for their valuable comments. This research was supported in part by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research [Nos. 23K22522 and 23H04899 (R. Yamazaki), 21H01082 and 23K20845 (N. Sago), 23H01176, 23K25872 and 23H04520 (H. Takahashi)] and Grant-in-Aid for JSPS Fellows [No. 22KF0329 (M. Meyer-Conde)]. This research was supported by the Joint Research Program of the Institute for Cosmic Ray Research, University of Tokyo, and Tokyo City University Prioritized Studies.
@@ A.1 Einstein Telescope
The Einstein Telescope (ET) [30] is an advanced third-generation GW detector proposed for construction underground in Europe. While its optical layout has not yet been finalized, various configurations have been proposed [60]. The proposed configurations are based on an equilateral triangle arrangement of three detectors, each consisting of two interferometers with 10-km long arms. One of the two interferometers is optimized for low-frequency bands utilizing low laser power and cryogenic mirrors. In contrast, the other is optimized for high-frequency bands utilizing high laser power and mirrors at room temperature. We use the ‘ET-D’ sensitivity curve [61] which is provided by PyCBC software [62]. We use this sensitivity curve divided by $\sin(60^{\circ})$ to make it for a triangular interferometer because the original sensitivity is based on simulations for an L-shaped interferometer.
@@ Appendix B Derivation of formulas
Considering the situation of accelerating jet, the Lorentz factor has a relation as follows:
@@
The acceleration time of the jet $t_{\mathrm{acc}}$ is written as:
@@
The geometric representation of the observer time is shown in Fig. 17, where $d$ is the distance from the projected point to the observer, which point is along with an arc from the position of $r_{0}$ at angle $\theta$ to the line of a hypothetical point particle. The observer start time $T_{s}(\theta)$ is expressed as:
@@
and observer end time $T_{E}(\theta,\gamma)$ is,
@@
The observer time $T_{\mathrm{obs}}(\theta,\gamma)$ is obtained by,
@@
and $\delta t\sim W/c$ is the expression under the ultra-relativistic speed.
Suppose that there are $n$ thin shells, where the thin shells are equally distributed, and the mass is rewritten as $m=nm^{\prime}$. We take the limit $n\to\infty$ in Eq. (12) by keeping the energy $\gamma nm^{\prime}c^{2}$ as constant. For simplicity, only term that depends on the number of shells $k$ are considered, and the expression is as follows:

### v2

@@ IV.2 Nearby GRB 221009A-like event
An energetic event GRB 221009A occurred nearby (redshift $z=0.151$, corresponding to the luminosity distance of 724 Mpc) [39, 40]. Its bright observed flux in the prompt GRB emission phase reaches Earth every $\sim 10^{4}$ yr. The isotropic equivalent gamma-ray energy of this event was about $E_{{\rm iso,}\gamma}\simeq 1.2\times 10^{55}$ erg. Well-sampled afterglows, spanning from radio to very-high-energy gamma rays, suggest that the event was associated with a narrowly collimated jet characterized by an opening half angle of $\Delta\theta\sim 0.01$ rad or smaller [41, 42, 43, 44, 45, 46, 47]. Then, the collimation-corrected energy of the gamma-ray emission is estimated as $(\Delta\theta)^{2}E_{{\rm iso,}\gamma}/4\simeq 3\times 10^{50}$ erg. If we assume the efficiency of the prompt GRB emission of 10%, the jet kinetic energy may be $\mathcal{E}\sim 3\times 10^{51}~\mathrm{erg}$.
@@
If similar events to GRB 221009A occurred very nearby, then the GW memory is detectable. Because the light curve of GRB 221009A consists of roughly five pulses in $\sim 510$ seconds [40], we adopt $N=5$ and total time $T_{\mathrm{CE}}\sim 510$ s. The distance to the source is assumed to be approximately one twentieth that of GRB 221009A, i.e., $R=30$ Mpc.
@@
The waveform is plotted in Fig. 14. It shows the detectability to observe the GW memory around 10 Hz and 100 Hz by ET-D. The attenuation at low frequency begins from $\sim 10^{-3}$ Hz because $fT_{\mathrm{CE}}\geq 1$ with $T_{\mathrm{CE}}^{-1}\sim 10^{-3}$ Hz, which are the exponents of exponential function in Eq. (13). The attenuation due to the shell thickness should appear at the frequency $\delta t_{j}^{-1}\sim(W_{j}/c)^{-1}\sim 10^{4}$ Hz; however, those amplitudes are already below the sensitivity curve of ET-D. This is because the attenuation associated with $T_{\mathrm{obs}}^{-1}\sim 10^{2}$~Hz begins earlier than that of $\delta t_{j}$ under the conditions $\gamma\in[300,600]$, $r_{0}=10^{7}~\mathrm{cm}$, and $\theta\in[0,0.4]$. The characteristic strain determined by Eq. (16) at low frequency is $\sim 10^{-23}$, which is observed around $10^{-4}$ Hz in Fig. 14. Furthermore, because the magnitude of characteristic strain is almost the same as that observed in the ET-D around $10^{2}$ Hz and is determined by $R$ and $\mathcal{E}$, it may contribute to the estimation of $\mathcal{E}$ of the GRBs with only ET-D observations.
@@
For the parameters given above, we find that the maximum detectable distance to the source is found to be approximately $\sim 20$ Mpc for ET-D and $\sim 60$ Mpc for DECIGO. If the distance is about 20 Mpc, our hypothetical event is simultaneously detectable by both detectors. Assuming that GW memory from GRB 221009A-like jets is detectable up to a distance of 60 Mpc, the likelihood of observing an on axis jet is extremely low: about one event every $\sim 2\times 10^{7}$ yrs. However, owing to the anti-beaming effect of GW memory emission, the jet’s GW signal is detectable even for off axis viewing angles. With a beaming factor of $f_{b}=2\pi/\Delta\Omega\sim 2\times 10^{4}$, the detection rate is boosted to roughly one event every $\sim 10^{3}$ yrs. This estimate neglects the dependence of the signal amplitude on the viewing angle $\theta_{v}$. Furthermore, the opening angle of the core of the jet of GRB 221009A is currently still uncertain. Hence our estimation on the rate would be greatly changed [41, 42, 43, 44, 45, 46, 47]. Although the detection rate of such events may remain low, nearby events with GW memory could provide valuable insights into the jet launching mechanism.
Let us suppose a more energetic event at a larger distance, such as $\mathcal{E}\sim 10^{52}$ erg and $R=500$ Mpc. The resulting waveforms for various combinations of angles $\theta_{v}$ and $\Delta\theta$ are shown in Fig. 15. Even at such a large distance, the GW memory is still detectable by DECIGO when observed off-axis, i.e., in the case of $\theta_{v}=10\times\Delta\theta$, which is roughly consistent with the result given in Ref. [7]. Note that in this paper, the effects of cosmic expansion are not taken into account. However, the luminosity distance of 500 Mpc corresponds to a redshift $z\sim 0.1$. Hence, they do not still qualitatively alter our argument.
@@
Figure 15: The waveform, using the same parameters as in Fig. 14, except for the distance to the source $R$, the total energy $\mathcal{E}$, and the combination of angles $\theta_{v}$ and $\Delta\theta$.
@@ IV.3 Possible GW emission from a black hole formation event in M31
Recently, the disappearance of a massive star in M31 was observed, marking the formation of a black hole [48]. The progenitor star, with an initial mass of approximately $20M_{\odot}$, evolved into a red supergiant before collapsing. At the terminal nuclear burning phase, its mass was estimated at $6.7M_{\odot}$ , and the star’s core collapsed and turned into a black hole with a mass $6.5M_{\odot}$. During its core collapse, a part of the stellar material was ejected. One can expect that some of it will likely generate a relativistic jet. However, the lack of bright optical emission suggested that, even if such a jet was launched, it likely did not escape from the star. If the GW memory had been detected, we would have obtained evidence of the jet in the star.
We simulated GW memory from M31 and discussed its detectability. The distance to the source M31 is relatively close, $R=0.77$ Mpc, and we typically set the parameters such as $T_{\mathrm{CE}}\sim 100$ sec and $\gamma_{j}\in[100,300]$. The hypothetical GW memory from M31 based on this event is shown in Fig. 16. Regarding the solid cyan line, we see that there is the initial attenuation at $10^{-2}$ Hz, induced by $T_{\mathrm{CE}}^{-1}\sim 10^{-2}$ Hz, and further attenuation around $T_{\mathrm{obs}}^{-1}\sim 10^{3}$ Hz. The thickness impact appears around $\delta t_{j}^{-1}\sim 10^{3}$ Hz, although these amplitudes are under the sensitivity curve of ET-D. The maximum amplitude given by Eq. (16) is $\sim 10^{-23}$, which is almost the same as the signal amplitude in the frequency range observable by ET-D. This indicates that a GW memory observed by ET-D is helpful to understand the total energy $\mathcal{E}$ of GRBs, which is the same situation in Fig. 14. Furthermore, the detection by DECIGO and ET-D depends on the $\mathcal{E}$. The dashed purple line in Fig. 16 shows the waveform under the same conditions as the solid cyan line, except for a smaller $\mathcal{E}$. This indicates that even for nearby GW events, detection remains challenging with next-generation detectors when the energy of the central engine is relatively small, around $10^{50}~\mathrm{erg}$.
@@ V Summary and Discussion
The time interval of each shell $\delta\tau_{j}$ with multiple shells affects the waveform shape. When the $\delta\tau_{j}={\rm const}$, which means $W_{j}={\rm const}$, the waveform has multiple peaks at specific frequencies due to the vanishing attenuation. If such GW memories are observed, the shells must be emitted periodically. In contrast, the waveform loses such specific peaks when the $\delta\tau_{j}$ exhibits disorder, such as each shell emitted on a different interval time $\delta\tau_{j}$.
@@
Furthermore, we have shown that the GW memories from relatively energetic phenomena with typical GRB parameters would be detectable by multiple GW detectors, ET-D and DECIGO.
After the initial acceleration, the GW amplitude on the low-frequency side is saturated, which is determined by the total energy in the jet. In this sense, the GW memory can serve as a form of calorimetry. However, after the initial acceleration, the jet evolves as it propagates as introduced in Sec. I. The amplitude of the GW memory varies with time, and its detailed analysis may tell us the total jet energy. Once the jet energy is known, we can estimate the efficiency of the prompt gamma-ray emission, whose constraints on the emission mechanism are currently unknown. Note that the order of magnitude of the jet kinetic energy of low-luminosity GRBs has not yet been rigorously determined. A popular model is in the low energy jet, while the off axis jet model gives larger energy [49, 50, 51, 52].
This paper shows that the GW memory from accelerating jets is only detectable for nearby sources. However, multimessenger signals other than GW are likely to be expected, if the GW memory signatures are detected. In most cases, the jet may be viewed offaxis. Then, electromagnetic emission is dim in gamma rays, and typical internal shock emission may be seen in the x-ray or ultraviolet/optical band due to the relativistic beaming effect [20, 53, 54]. The prompt GRB emission may not be detected, and only off axis afterglow emission may be observable [55, 56, 57]. If the GRB jet contains baryons to some extent, one can also expect high-energy (TeV–PeV) neutrinos [58, 37, 59, 60, 61, 62]. When the jet is viewed off axis, the neutrino emission is again dim but may still be observable [63].
@@ Acknowledgements.
We would like to thank Drs. K. Ioka, S. J. Tanaka, S. Kimura and K. Komori for their valuable comments. This research was supported in part by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research [No. 23K22522 and No. 23H04899 (R. Y.): No. 21H01082 and No. 23K20845 (N. S.): No. 23H01176, 23K25872 and No. 23H04520 (H. T.)] and Grant-in-Aid for JSPS Fellows [No. 22KF0329 (M. M.-C.)]. This research was supported by the Joint Research Program of the Institute for Cosmic Ray Research, University of Tokyo, and Tokyo City University Prioritized Studies.
Data Availability
The data that support the findings of this article are openly available [64]; embargo periods may apply.
@@ A.1 Einstein Telescope
The ET [31] is an advanced third-generation GW detector proposed for construction underground in Europe. While its optical layout has not yet been finalized, various configurations have been proposed [65]. The proposed configurations are based on an equilateral triangle arrangement of three detectors, each consisting of two interferometers with 10-km long arms. One of the two interferometers is optimized for low-frequency bands utilizing low laser power and cryogenic mirrors. In contrast, the other is optimized for high-frequency bands utilizing high laser power and mirrors at room temperature. We use the “ET-D” sensitivity curve [66] which is provided by PyCBC software [67]. We use this sensitivity curve divided by $\sin(60^{\circ})$ to make it for a triangular interferometer because the original sensitivity is based on simulations for an L-shaped interferometer.
@@ Appendix B DERIVATION OF FORMULAS
Considering the situation of the accelerating jet, the Lorentz factor has a relation as follows:
@@
The acceleration time of the jet $t_{\mathrm{acc}}$ is written as
@@
The geometric representation of the observer time is shown in Fig. 18, where $d$ is the distance from the projected point to the observer, the point being along with an arc from the position of $r_{0}$ at angle $\theta$ to the line of a hypothetical point particle. The observer start time $T_{s}(\theta)$ is expressed as
@@
and the observer end time $T_{E}(\theta,\gamma)$ is
@@
The observer time $T_{\mathrm{obs}}(\theta,\gamma)$ is obtained by
@@
and $\delta t\sim W/c$ is the expression under the ultrarelativistic speed.
Suppose that there are $n$ thin shells, where the thin shells are equally distributed, and the mass is rewritten as $m=nm^{\prime}$. We take the limit $n\to\infty$ in Eq. (12) by keeping the energy $\gamma nm^{\prime}c^{2}$ as constant. For simplicity, only terms that depend on the number of shells $k$ are considered, and the expression is as follows: